For the following reaction, 16.2 grams of chlorine gas are allowed to react with 49.6 grams of sodium iodide chlorine(g sodium iodide(s sodium chloride(s) iodine(s) What is the maximum amount of sodium chloride that can be formed? grams What is the FORMULA for the limiting reagent? What amount of the excess reagent remains after the reaction is complete? grams
Answer #1

Cl2 (g) + 2 NaI (s)   ----> 2 NaCl (s) + I2 (s)

mol of Cl2 available = mass of Cl2 / molar mass of Cl2
= 16.2/71
=0.228 mol

mol of NaI available = mass of NaI / molar mass of NaI
= 49.6/150
=0.331 mol

According to balanced reaction, mol of NaI required is twice the moles of Cl2
But NaI available is less
So, NaI is limiting reagant

a)
from reaction,
mol of NaCl formed = moles of NaI = 0.331 mol
mass of NaCl = number of moles * molar mass
= 0.331 mol * 58.5 g/mol
= 19.4 g
Answer: 19.4 g

b)
NaI is limiting reagant

c)
moles of Cl2 reacted = (1/2)*moles of NaI reacted
= (1/2)*0.331
= 0.165 mol

mol of Cl2 remaining = 0.228 - 0.165 = 0.063 mol
mass of Cl2 remaining = mol remaining * molar mass of Cl2
= 0.063 mol * 71 g/mol
= 4.47 g

Answer: 4.47 g